OpenStudy (anonymous):

125% as a fraction please and the dabbleboard web sight no one is on line to help.

OpenStudy (anonymous):

I'm online there just type in the chat box http://www.dabbleboard.com/draw?b=Guest663757&i=0&c=eb82ae69c0230a71d5a9894db588a7bbb2e5f7bf

OpenStudy (anonymous):

125/100=1.25=5/4

OpenStudy (anonymous):

0.125 = 125/1000 = 1/8

OpenStudy (anonymous):

too quick on the fingers...I apologize the above is correct

OpenStudy (anonymous):

but in the answer key it says 1 1 over 4 ?

OpenStudy (anonymous):

I think your answer key is wack! lol every fraction is over 100. so put 125 over 100 and simplify

OpenStudy (anonymous):

125/100 you can simplfy this by dividing both the top and bottom by 25 125/25=5 100/25=4 5/4 is your answer

OpenStudy (anonymous):

thanks all of u

OpenStudy (anonymous):

no problem di you figure out the 3/4 one?

OpenStudy (anonymous):

125 over 100 = 5 over 4 simplest form 1 1 over 4

OpenStudy (anonymous):

got it thanks

OpenStudy (anonymous):

ok good